# A Single Online Agent Can Efficiently Learn Mean Field Games

### v1

@@ Abstract
mean field games, online learning, model-free, single-agent
@@ 1 Introduction
  • We prove that QMI efficiently approximates FPI and, therefore enjoys a similar convergence guarantee. The resemblance between the learning dynamics of QMI and FPI is illustrated in Figure 2. We provide sample complexity guarantees for our methods (Theorem 1). Our methods are the first provably efficient online single-agent model-free methods for learning MFGs. We validate our findings through numerical experiments on various MFGs (LABEL:sec:exp and LABEL:sec:apx-exp).
@@ 3 Online Stochastic Updates
where $\beta$ is the step size. The full derivation of this update rule is deferred to LABEL:sec:apx-update. Similar to TD control, we anticipate that this update rule drives the M-value function to converge to the population distribution induced by $P_{\scriptscriptstyle Q}$. Furthermore, selecting a step size of $\beta_{t}=1/(t+1)$ simplifies it to a Markov chain Monte Carlo (MCMC) method, validating its correctness.
@@
Assumption 2 is a standard assumption for online RL methods [bhandari2018FiniteTime, zou2019Finitesampleanalysis, qu2020Finitetimeanalysis].
@@ 5 Sample Complexity Analysis
return the updated Q- and M-value function after an outer iteration of Algorithm 1, consisting of $T$ online stochastic updates using 8 and 9. Then, the off-policy QMI operator is the composition of the above two operators: $\hat{\Gamma}_{\mathrm{off}}(T)\coloneqq\Gamma_{M}(T)\circ\Gamma_{Q}(T).$
@@
returns the updated M-value function after an outer iteration of Algorithm 1, consisting of $T$ online stochastic updates using 8 and 9, as well as the policy updates using 13.
@@
(a) Mean squared error
(b) Exploitability
(c) Learned population distributions
Figure 3: Performance comparison of FPI, off-policy QMI, and on-policy QMI on ring road speed control. MSE represents the mean squared $L_{2}$ error between the current M-value function and the MFNE population distribution. Exploitability refers to the disparity between the current value function and the BR w.r.t. the current population distribution (see LABEL:sec:apx-exp and perrin2020Fictitiousplay). Learned population distributions are scaled for better visualization.
(a) Mean squared error
(b) Exploitability
(c) Learned population distribution
Figure 4: Performance comparison of FPI, off-policy QMI, and on-policy QMI on Sioux Falls network routing. Only the population distribution learned by off-policy QMI is shown in (c); other methodsgive similar population distributions and are deferred to LABEL:sec:exp-rr.
@@
where the backtracking discrepancy $H_{1}$ and the distribution progress $H_{3}$ are controlled by the virtual period $\tau$ (LABEL:lem:progress and LABEL:lem:back), while the two convergence rates $H_{2}$ and $H_{4}$ are characterized by the geometric ergodicity of stationary MDPs and the sample complexity of SARSA (LABEL:lem:mix and LABEL:lem:control), respectively. In brief, we show that the agent’s state distribution, and thus its M-value function, rapidly converges to the IP $\mu_{M}$.
@@
with the number of iterations being at most

### v2

@@ 1 Introduction
  • We prove that QMI efficiently approximates FPI and, therefore enjoys a similar convergence guarantee. The resemblance between the learning dynamics of QMI and FPI is illustrated in Figure 2. We provide sample complexity guarantees for our methods (Theorem 1). Our methods are the first provably efficient online single-agent model-free methods for learning MFGs. We validate our findings through numerical experiments on various MFGs (LABEL:sec:exp).
@@ 3 Online Stochastic Updates
where $\beta$ is the step size. Please refer to LABEL:sec:apx-update for the full derivation of this update rule. Similar to TD control, we anticipate that this update rule drives the M-value function to converge to the population distribution induced by $P_{\scriptscriptstyle Q}$. Furthermore, selecting a step size of $\beta_{t}=1/(t+1)$ simplifies it to a Markov chain Monte Carlo (MCMC) method, validating its correctness.
@@
Assumption 2 is a standard assumption for online methods [bhandari2018FiniteTime, zou2019Finitesampleanalysis, qu2020Finitetimeanalysis].
@@ 5 Sample Complexity Analysis
return the updated Q- and M-value function after an outer iteration of Algorithm 1, consisting of $T$ online stochastic updates using Lines 8 and 9. Then, the off-policy QMI operator is the composition of the above two operators: $\hat{\Gamma}_{\mathrm{off}}(T)\coloneqq\Gamma_{M}(T)\circ\Gamma_{Q}(T).$
@@
returns the updated M-value function after an outer iteration of Algorithm 1, consisting of $T$ online stochastic updates using Lines 8 and 9, as well as the policy updates using Line 13.
@@
((a)) Mean squared error
((b)) Exploitability
((c)) Learned population distributions
Figure 3: Performance comparison of FPI, off-policy QMI, and on-policy QMI on ring road speed control. MSE represents the mean squared $L_{2}$ error between the current M-value function and the MFNE population distribution. Exploitability refers to the disparity between the current value function and the BR w.r.t. the current population distribution. Learned population distributions are scaled for better visualization.
((a)) Mean squared error
((b)) Exploitability
((c)) Learned population distribution
Figure 4: Performance comparison of FPI, off-policy QMI, and on-policy QMI on Sioux Falls network routing. Only the population distribution learned by off-policy QMI is shown in (c); other methodsgive similar population distributions (please refer to LABEL:sec:exp-rr).
@@
where the backtracking discrepancy $H_{1}$ and the distribution progress $H_{3}$ are controlled by the virtual period $\tau$ (LABEL:lem:progress and LABEL:lem:back), while the two convergence rates $H_{2}$ and $H_{4}$ are characterized by the geometric ergodicity of stationary MDPs and the sample complexity of SARSA (LABEL:lem:control and LABEL:lem:mix), respectively. In brief, we show that the agent’s state distribution, and thus its M-value function, rapidly converges to the IP $\mu_{M}$.
@@
where $\hat{\Gamma}$ can be either $\hat{\Gamma}_{\mathrm{off}}$ or $\hat{\Gamma}_{\mathrm{on}}$, with the number of iterations being at most